RC Meeting, Monday, August 20

From HallCWiki
Jump to: navigation, search
  • General
    • No hall lockup tonight
    • Beam is possible Tuesday night, but not likely. Will not cancel shifts until certain though.
    • Issues with the upstream beamline vacuum. Ion pumps not yet on
    • Still work to be done in hall before ready for lockup
    • Need to look into merging the KAON-LT shift instruction with existing instructions on the wiki
    • New procedure for "on call" - will need to update list on wiki to reflect new on-call list
  • Target
    • Scattering chamber leak repaired over the weekend (big thanks to Dave Meekins, Chris Keith, and Walter Kellner for coming in)
    • Target cooldown underway
  • SHMS
    • HGC at 1 atm of C4F8O
    • No bubbles at exit of DC1, but gas seems to be flowing
  • HMS
    • Cerenkov window repaired and installed
    • Cerenkov at 0.45 atm, C4F8O
  • DAQ
    • Hall B will be running in "delayed helicity reporting" - likely not an issue (for now)